c) What are the dimensions of heat capacity? Does this make sense? Explain.
Answer #1

We know that,

Heat = m * heat capacity * temperature difference

Dimensions of Heat = ML^2T^-2

where M is dimension of mass, L is for length and T is for time.

Dimension of temperature = K

Now,

ML^2T^-2 = M * s * K

Dimension of heat capacity = L^2T^-2K^-1